Hat sizes are determined by measuring the circumference of one's head in either inches or centimeters. Use ratio and proportion to complete the following table.
Hat size Head Circumference Head Circumference
(to nearest 1/5 inch) (to nearest centimeter)

6 7/8 21 3/5 55
? ? 62

The first is rounded to the nearest eith inch and second is rounded to neares fith inch.

Answer by stanbon(75887)

Hat sizes are determined by measuring the circumference of one's head in either inches or centimeters. Use ratio and proportion to complete the following table.
Hat size Head Circumference Head Circumference
(to nearest 1/5 inch) (to nearest centimeter)

6 7/8 21 3/5 55
? ? 62

The first is rounded to the nearest eith inch and second is rounded to neares fith inch.
----------------
x/(6 7/8) = 62/55
---
x/(55/8) = 62/55
---
8x/55 = 62/55
8x = 62
x = 31/4 = 7 3/4
-----------------------------
x/(21 3/5) = 62/55
x/(108/5) = 62/55
5x/108 = 62/55
5x = 108(62/55)
x = 24.3491
==============================
Cheers,
Stan H.